OpenStudy (anonymous):

this one is annoying; 15 - g = 23 - 2g step by step would be great, thanks!

OpenStudy (anonymous):

15-g=23-2g 15+g=23+g 15=23+3g -15=-15 8=3g 8/3=3g/3 G=22/8 G=2 1/4 <- Answer Hope I helped give me a medal if I did PLEASE, if I helped after u give me a reward please close questions.... ANy questions further become my fan, then message me with ur link please

OpenStudy (anonymous):

15-g=23-2g to isolate the g term by itself add the g on the left to the g on the right +g +g 15=23 -1g now to get the g term by itself subtract the 23 to the 15 -23 -23 -8=22g then divide 22 to -8 to get your final answer. -8/22=g

OpenStudy (anonymous):

hope I helped?????

OpenStudy (cynthiers):

15-g=23-2g +g +g 15=23-g g=23 -15 g=8

OpenStudy (anonymous):

thankyou guys